OpenStudy (anonymous):

Part II: Dividing Powers to Solve Real-world Problems Set up and simplify a division of powers expression to solve each problem. Give your answer in both exponential form and as a whole number, showing all work leading to your answer. Then write a sentence to explain your answer. 9. The population of the African country Uganda has been growing at a steady rate for several decades. Assuming that the current population is about 225 million and that the country has an approximate land area of 217 square miles, about how many people per square mile live in Uganda

OpenStudy (anonymous):

help please :(

OpenStudy (anonymous):

please help...?

OpenStudy (whpalmer4):

population is 225 million. What's 1 million written as a power of 10? The area of Uganda is much more than 217 square miles!

OpenStudy (whpalmer4):

and the population is much less than 225 million. Let's hear it for bogus questions...

OpenStudy (whpalmer4):

Well, let's say it was a population of 319 million, and an area of 3.719 million square miles. You can write 1 million as \(10^6\), so make the population \(319*10^6\) people. That area would be \(3.719*10^6\) square miles. Population density would be people/area = \[\frac{319*10^6 \text{ people}}{3.719*10^6 \text{ mi}^2} = \frac{319\text{ people}}{3.719\text{ mi}^2} = 85.77\text{ people/mi}^2 \]